OpenStudy (anonymous):

Rearrange this equation to isolate C a=b(1/c-1/d)

OpenStudy (anonymous):

Divide both sides by b: a/b = 1/c - 1/d add 1/d to both sides a/b + 1/d = 1/c combine the two fractions on the left into one note that if x/y = 1/c then c = y/x and use this principle on what you got above.

OpenStudy (anonymous):

I understand everything except how to combine the fractions on the left.

OpenStudy (anonymous):

a/b + 1/d = da/db + b/db = (da + b)/db so they have same denominator next flip this and 1/c

OpenStudy (anonymous):

so c =bd/ad+b ?